OpenStudy (anonymous):

A trapezoid ABCD is drawn on a coordinate grid as shown below. Square ABCD. Each vertex labeled A(1, -2) B(1, 1) C(4, 1) D(6, -2) Sean will draw a trapezoid EFGH similar to trapezoid ABCD. The length of EF should be 6 and point E should be located at (0, -2). At which points will vertices F and H be located? F(0, 4), H(10, -2) F(0, 4), H(6, 4) F(7, 7), H(12, 4) F(6, 4), H(2, -2)

OpenStudy (anonymous):

thats the image

OpenStudy (anonymous):

How do u spell ur name Machi303?

OpenStudy (anonymous):

mach 1303

OpenStudy (anonymous):

Mach as in Mac or Mak??

OpenStudy (anonymous):

michelleine i add my first and last name together to get that

OpenStudy (anonymous):

thats my name michelleine

OpenStudy (jdoe0001):

hmm, fyi flvs urls are usually session localized resources meaning unless you're logged in flvs website, you can't access the resource take a screenshot of the material and post it here

OpenStudy (anonymous):

if you click on it it give u the image

OpenStudy (jdoe0001):

yes it did, I guess this resource isn't localized then that isn't the norm btw with flvs

OpenStudy (anonymous):

yeah can u help me

OpenStudy (jdoe0001):

well the length of EF will be 6 units notice the vertex are being listed clockwise, starting from A ... D the new SIMILAR trapezoid is listed as EFGH if we start clockwise also from point E, it'd look like  so the EF side on the bigger trapezoid, will correspond with the AB line on the smaller trapezoid now that we know the correspondence of the sides we can also just redraw EFGH facing the other side like 

OpenStudy (jdoe0001):

so line EF will be = 6 how long is its corresponding side on the smaller trapezoid of AB, that is, how long is AB?

OpenStudy (anonymous):

so what would be the vertices i don't get it

OpenStudy (jdoe0001):

well, we dunno the vertices yet but since they're SIMILAR polygons, you want to know their ratio of correspondence

OpenStudy (jdoe0001):

how long is AB?

OpenStudy (anonymous):

one sec

OpenStudy (anonymous):

3units

OpenStudy (jdoe0001):

so AB is 3 EF is 6 both sides are correspondent of each other so EF is really twice as big as AB 6/3 = 2 so the ratio of proportion between both SIMILAR polygons is 2 to 1 so if you want to get any side length on EFGH just look at the corresponding side on ABCD and double that amount E is located at (0, -2) F is 6 units up from there AD on the smaller trapezoid is 5 units long so EH will be 5*2 = 10 E is located at (0, -2) H is located 10 units to the left from there

OpenStudy (anonymous):

i am supposed to find f

OpenStudy (jdoe0001):

yes E is located at (0, -2) F is 6 units up from there

OpenStudy (anonymous):

so it would be A

OpenStudy (jdoe0001):

(0, -2) 6 units up, that is (0, -2+6) = (0, 4) => F (0, -2) 10 units to the left, (0+10,-2) = (10, -2) => H

OpenStudy (jdoe0001):

yes

OpenStudy (anonymous):

thnx

OpenStudy (jdoe0001):

yw